OpenStudy (anonymous):

Fill in the missing # for the solution to this system of equations: 2x + 4y = 8 3x + 2y = 10 x = ?

OpenStudy (anonymous):

2x+4y=8 -6x-4y=-20 -4x=-12 x=3 i think